Problem 10-18 Return on Investment (ROI) and Residual Income (LO10-1, LO10-2] I know headquarters wants us to add that new pRequired: 1. Compute the Office Products Divisions ROI for this year. 2. Compute the Office Products Divisions ROI for the
Answer #1

1.ROI = Net Operating Income/Average Assets

= 800,000/4,000,000

= 20%

2.ROI of new product line = (2,000,000*40%-640,000)/1,000,000

= 16%

3.ROI when new line is added

= (800,000+160,000)/5,000,000

= 19.2%

4.Reject, since will reduce ROI

5.Because ROI of new line is higher than the overall ROI of the company

4.Residual Income = Net operating Income – Average operating assets*Minimum required rate of return

a.Office Products residual income = 800,000 – 4,000,000*12%

= $320,000

b.New Product Line added = 160,000 – 1,000,000*12%

= $40,000

c.Residual Income = 320,000 + 40,000 = $360,000

d.Accept as Residual Income has increased